A restaurant is owned by three partners. Partner a owns 3/5 of it, partner B owns twice as much as partner c. What percentage of the restaurant belongs to partner B?

Answer by richwmiller(17219)

a=3/5 or 60%
b+c=2/5
b=2c
b+c=2/5= 6/15 or 40%
b=4/15 and c=2/15
4/15 belongs to B
A owns 60%
B owns 26.67%
c owns 13.33%